# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### v4

@@ 5.2 Comparison with State-of-the-art
| STTR | 49.72 | 6.80 | 40.26 | 6.16 | OOM | 38.89 |
| PSMNet | 30.51 | 4.68 | 32.15 | 5.99 | 33.53 | 18.02 |
| CFNet | 13.64 | 2.27 | 12.09 | 2.89 | 23.91 | 7.67 |
| AANet | 7.23 | 1.27 | 7.72 | 1.41 | 22.45 | 18.77 |
| Mobilestereo-2D | 18.34 | 2.45 | 21.21 | 2.78 | 34.04 | 13.89 |
| Mobilestereo-3D | 18.96 | 2.79 | 19.69 | 3.40 | 29.32 | 13.71 |
| GwcNet | 23.05 | 2.76 | 25.19 | 3.58 | 29.87 | 14.54 |
| COEX | 12.08 | 1.80 | 11.00 | 2.48 | 25.17 | 11.43 |
| FADNet++ | 11.31 | 1.77 | 13.23 | 2.97 | 24.17 | 25.53 |
| CascadeStereo | 11.83 | 1.83 | 12.03 | 2.69 | 27.27 | 11.68 |
| IGEV | 7.80 | 1.40 | 6.35 | 1.31 | 10.51 | 4.44 |
| StereoBase | 4.85 | 0.99 | 5.35 | 1.18 | 9.76 | 3.12 |

### v5

@@ 5.2 Comparison with State-of-the-art
| STTR† | 49.72 | 6.80 | 40.26 | 6.16 | OOM | 38.89 |
| PSMNet† | 30.51 | 4.68 | 32.15 | 5.99 | 33.53 | 18.02 |
| CFNet† | 13.64 | 2.27 | 12.09 | 2.89 | 23.91 | 7.67 |
| AANet† | 7.23 | 1.27 | 7.72 | 1.41 | 22.45 | 18.77 |
| Mobilestereo-2D† | 18.34 | 2.45 | 21.21 | 2.78 | 34.04 | 13.89 |
| Mobilestereo-3D† | 18.96 | 2.79 | 19.69 | 3.40 | 29.32 | 13.71 |
| GwcNet† | 23.05 | 2.76 | 25.19 | 3.58 | 29.87 | 14.54 |
| COEX† | 12.08 | 1.80 | 11.00 | 2.48 | 25.17 | 11.43 |
| FADNet++† | 11.31 | 1.77 | 13.23 | 2.97 | 24.17 | 25.53 |
| CascadeStereo† | 11.83 | 1.83 | 12.03 | 2.69 | 27.27 | 11.68 |
| IGEV† | 7.80 | 1.40 | 6.35 | 1.31 | 10.51 | 4.44 |
| StereoBase† | 4.85 | 0.99 | 5.35 | 1.18 | 9.76 | 3.12 |